# Interpolating many-body wave functions for accelerated molecular dynamics on near-exact electronic surfaces

### v1

@@ Acknowledgments
We thank Huanchen Zhai for support with the block2 code, Hugh Burton for support with the libgnme code, Carlos Mejuto-Zaera for helpful insights about an independently-developed related scheme, Oliver Backhouse for technical help and Kemal Atalar and Lachlan Lindoy for additional feedback on the manuscript. We gratefully acknowledge support from the Air Force Office of Scientific Research under award number FA8655-22-1-7011 and the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1). YR also acknowledges the support of the Engineering and Physical Sciences Research Council [grant EP/Y005090/1].
@@
YR and GHB jointly developed the methodology and wrote the manuscript. YR implemented the approach and performed the numerical experiments.
@@
    where $\boldsymbol{\psi}$ is the inferred ground-state eigenfunction from the diagonalization of the Hamiltonian in the many-body basis of the fixed training states (Eq. (2) of the main text). Substituting in the definition of the Hamiltonian from Eq. (5), similar to the derivation in Ref. [10.1063/1.4927594], we obtain
@@
    The evaluation of the the one- and two-electron integral gradients in the SAO basis requires the derivative of the transformation $\mathbf{Z}(\mathbf{R})$ from the atomic orbitals to the SAO basis, as well as the gradients of the integrals between atomic orbitals. For the one-electron integral derivative, we obtain

### v2

@@ Acknowledgments
We thank Huanchen Zhai for support with the block2 code, Hugh Burton for support with the libgnme code, Carlos Mejuto-Zaera for helpful insights about an independently-developed related scheme, Oliver Backhouse for technical help and Kemal Atalar, Venkat Kapil and Lachlan Lindoy for additional feedback on the manuscript. We gratefully acknowledge support from the Air Force Office of Scientific Research under award number FA8655-22-1-7011 and the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1). YR also acknowledges the support of the Engineering and Physical Sciences Research Council [grant EP/Y005090/1].
@@
YR and GHB jointly developed the methodology and wrote the manuscript. YR implemented the approach and performed the numerical experiments. GHB supervised the project.
@@
where $\boldsymbol{\psi}$ is the inferred ground-state eigenfunction from the diagonalization of the Hamiltonian in the many-body basis of the fixed training states (Eq. 2 of the main text). Substituting in the definition of the Hamiltonian from Eq. 5, similar to the derivation in Ref. [10.1063/1.4927594], we obtain
@@
The evaluation of the the one- and two-electron integral gradients in the SAO basis requires the derivative of the transformation $\mathbf{Z}(\mathbf{R})$ from the atomic orbitals to the SAO basis (a Pulay-like force contribution), as well as the gradients of the integrals between atomic orbitals. For the one-electron integral derivative, we obtain